OpenStudy (anonymous):

WILL MEDAL AND FAN!!! The Rainier Furniture Factory takes twice as long to make a sofa as it does to make a chair. If the time to create 3 sofas and 4 chairs is no more than 15 hours, what is the longest time it could take to produce 2 sofas?

OpenStudy (anonymous):

As it takes twice to make a sofa as it does to make a chair, the time necessary to make two chairs equals the time of one sofa. So we go from: \[3s + 4c \le 15\] to: (as 2c = 1s) \[3s + 2s = 5s \le 15\] So, \[s \le 5\] Finally, as one sofa takes no more than 5 minutes, two sofas take twice as much: \[2s \le 10\] Two sofas take the longest time of 10 minutes.

OpenStudy (anonymous):

the problem says that I have to figure out the number of hours @RMoyses

OpenStudy (anonymous):

sorry I read minutes instead of hours, but it's exactly the same thing, just replace every "minute" I wrote for "hour". Hope you understand it =)

OpenStudy (anonymous):

Thanks so much @RMoyses